# SPEAR: Subset-sampled Performance Evaluation via Automated Ground Truth Generation for RAG

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) is a core approach for enhancing Large Language Models (LLMs), where the effectiveness of the retriever largely determines the overall response quality of RAG systems. Retrievers encompass a multitude of hyperparameters that significantly impact performance outcomes and demonstrate sensitivity to specific applications. Nevertheless, hyperparameter optimization entails prohibitively high computational expenses. Existing evaluation methods suffer from either prohibitive costs or disconnection from domain-specific scenarios. This paper proposes SPEAR (Subset-sampled Performance Evaluation via Automated Ground Truth Generation for RAG), which addresses evaluation data challenges through subset sampling techniques and achieves robust automated retriever evaluation by minimal retrieval facts extraction and comprehensive retrieval metrics. Based on real user queries, this method enables fully automated retriever evaluation at low cost, thereby obtaining optimal retriever for specific business scenarios. We validate our method across classic RAG applications in rednote, including knowledge-based Q&A system and retrieval-based travel assistant, successfully obtaining scenario-specific optimal retrievers.
SPEAR: Subset-sampled Performance Evaluation
via Automated Ground Truth Generation for RAG
@@ 1 Introduction
Therefore, we propose SPEAR, an automated approach for obtaining optimal retrievers that efficiently enables fully automated retriever evaluation using real user queries. The core advantages of our method include:
@@
This paper will introduce the details of the SPEAR method in the Core Challenges and Solutions section, employ this method in two RAG applications at rednote in the Cases section, and finally present other research work related to this paper in the Related Work section.
@@ 2.3 Cross-Retriever Evaluation
For instance, consider a comparison between retriever A using segmentation strategy A and retriever B using segmentation strategy B in Figure 2. For a specific query recalling a particular overlapping gt, retriever A recalls the segment "$redundancyA\|gt$", while retriever B recalls the segment "$gt\|redundancyB$". Both segments would be classified as relevant by the LLM and thus added to the GT subset, as both contain valid information for answering the question. However, during Pseudo GT generation, "$redundancyA\|gt$" and "$gt\|redundancyB$" would be counted as two distinct gt chunks, leading to the erroneous conclusion that retriever A failed to recall "$gt\|redundancyB$" and retriever B failed to recall "$redundancyA\|gt$". Such computational results are unreasonable, as both retriever A and retriever B have effectively recalled the essential gt.
@@
For example, in the case above, the LLM would segment the recalled "$redundancyA\|gt$", identifying "redundancy A" as irrelevant redundant text and extracting the minimal relevant text block gt. Similarly as recalled "$gt\|redundancyB$", thereby ensuring that both retriever A and retriever B successfully recall gt during precision and recall calculations, achieving universal relative evaluation across retrievers.
@@ 3 Cases
Given the current state of RAG system applications in industry, this paper selects two representative RAG application within rednote: the knowledge-based QA application represented by "Linghang Shu" and the generative retrieval application in travel scenarios. We employ the proposed SPEAR method to automatically obtain optimal retrievers for both application.
@@ 3.1.1 Retriever Construction Strategies
Our evaluation queries based on 266 real user queries obtained from online service. We employ the SPEAR method to obtain the optimal retriever.
@@ 5 Conclusion
Addressing the challenge of determining optimal retriever construction strategies for specific business scenarios in RAG systems, this paper proposes the SPEAR method. Through the utilization of real user queries, this method enables fully automated retriever evaluation, thereby identifying optimal retrievers for specific business contexts, providing a data-driven solution for RAG system performance improvement. Based on this method, we automatically obtain optimal retrievers in two scenarios in rednote. Experimental results demonstrate that optimal retrievers are highly application-dependent, the automated acquisition of optimal strategy combinations tailored to diverse application environments through the SPEAR method holds significant implications for the practical performance of RAG applications.
@@ 6 Limitation
Trade-off between computational efficiency and evaluation accuracy: Our proposed SPEAR method faces a challenge in balancing computational cost with evaluation quality. When constructing Pseudo GT through subset sampling, incorporating a larger number of retrievers in the comparison process yields Pseudo GT that more closely approximates the true GT, thereby enhancing evaluation reliability. However, this improved accuracy comes at the expense of significantly increased computational overhead, as each additional retriever requires substantial processing time and resources. This computational burden becomes particularly pronounced in large-scale retrieval systems or when frequent evaluations are necessary. From a practical standpoint, different application scenarios may require different evaluation strategies - real-time systems prioritize computational efficiency while offline evaluation systems can afford more comprehensive comparisons. Future work should focus on developing adaptive mechanisms that can dynamically adjust the evaluation scope based on available computational resources and required accuracy levels.
Dependence on authentic user queries and handling of unanswerable queries: SPEAR’s effectiveness is inherently tied to the availability and quality of authentic user query datasets, which may not always be accessible or representative. The method assumes that user queries have retrievable answers within the knowledge base, but real-world scenarios often include unanswerable queries - those for which no relevant information exists. Currently, our framework does not distinguish between answerable and unanswerable queries, which may lead to suboptimal evaluation outcomes and wasted computational resources. This limitation stems from the challenge of automatically detecting query answerability without access to ground truth annotations. Developing robust mechanisms to identify and appropriately handle unanswerable queries represents a promising direction for future work, potentially through query difficulty estimation or retrieval confidence scoring techniques.

### v2

# SEARA: An Automated Approach for Obtaining Optimal Retrievers

@@ Abstract
Retrieval-Augmented Generation (RAG) is a core approach for enhancing Large Language Models (LLMs), where the effectiveness of the retriever largely determines the overall response quality of RAG systems. Retrievers encompass a multitude of hyperparameters that significantly impact performance outcomes and demonstrate sensitivity to specific applications. Nevertheless, hyperparameter optimization entails prohibitively high computational expenses. Existing evaluation methods suffer from either prohibitive costs or disconnection from domain-specific scenarios. This paper proposes SEARA (Subset sampling Evaluation for Automatic Retriever Assessment), which addresses evaluation data challenges through subset sampling techniques and achieves robust automated retriever evaluation by minimal retrieval facts extraction and comprehensive retrieval metrics. Based on real user queries, this method enables fully automated retriever evaluation at low cost, thereby obtaining optimal retriever for specific business scenarios. We validate our method across classic RAG applications in rednote, including knowledge-based Q&A system and retrieval-based travel assistant, successfully obtaining scenario-specific optimal retrievers.
SEARA: An Automated Approach for Obtaining Optimal Retrievers
@@ 1 Introduction
Therefore, we propose SEARA, an automated approach for obtaining optimal retrievers that efficiently enables fully automated retriever evaluation using real user queries. The core advantages of our method include:
@@
This paper will introduce the details of the SEARA method in the Core Challenges and Solutions section, employ this method in two RAG applications at rednote in the Cases section, and finally present other research work related to this paper in the Related Work section.
@@ 2.3 Cross-Retriever Evaluation
For instance, consider a comparison between retriever A using segmentation strategy A and retriever B using segmentation strategy B in Figure 2. For a specific query recalling a particular overlapping gt, retriever A recalls the segment $redundancyA\|gt$", while retriever B recalls the segment $gt\|redundancyB$". Both segments would be classified as relevant by the LLM and thus added to the GT subset, as both contain valid information for answering the question. However, during Pseudo GT generation, $redundancyA\|gt$" and $gt\|redundancyB$" would be counted as two distinct gt chunks, leading to the erroneous conclusion that retriever A failed to recall $gt\|redundancyB$" and retriever B failed to recall $redundancyA\|gt$". Such computational results are unreasonable, as both retriever A and retriever B have effectively recalled the essential gt.
@@
For example, in the case above, the LLM would segment the recalled $redundancyA\|gt$", identifying "redundancy A" as irrelevant redundant text and extracting the minimal relevant text block gt. Similarly as recalled $gt\|redundancyB$", thereby ensuring that both retriever A and retriever B successfully recall gt during precision and recall calculations, achieving universal relative evaluation across retrievers.
@@ 3 Cases
Given the current state of RAG system applications in industry, this paper selects two representative RAG application within rednote: the knowledge-based QA application represented by "Linghang Shu" and the generative retrieval application in travel scenarios. We employ the proposed SEARA method to automatically obtain optimal retrievers for both application.
@@ 3.1.1 Retriever Construction Strategies
Our evaluation queries based on 266 real user queries obtained from online service. We employ the SEARA method to obtain the optimal retriever.
@@ 5 Conclusion
Addressing the challenge of determining optimal retriever construction strategies for specific business scenarios in RAG systems, this paper proposes the SEARA method. Through the utilization of real user queries, this method enables fully automated retriever evaluation, thereby identifying optimal retrievers for specific business contexts, providing a data-driven solution for RAG system performance improvement. Based on this method, we automatically obtain optimal retrievers in two scenarios in rednote. Experimental results demonstrate that optimal retrievers are highly application-dependent, the automated acquisition of optimal strategy combinations tailored to diverse application environments through the SEARA method holds significant implications for the practical performance of RAG applications.
@@ 6 Limitation
Trade-off between computational efficiency and evaluation accuracy: Our proposed SEARA method faces a challenge in balancing computational cost with evaluation quality. When constructing Pseudo GT through subset sampling, incorporating a larger number of retrievers in the comparison process yields Pseudo GT that more closely approximates the true GT, thereby enhancing evaluation reliability. However, this improved accuracy comes at the expense of significantly increased computational overhead, as each additional retriever requires substantial processing time and resources. This computational burden becomes particularly pronounced in large-scale retrieval systems or when frequent evaluations are necessary. From a practical standpoint, different application scenarios may require different evaluation strategies - real-time systems prioritize computational efficiency while offline evaluation systems can afford more comprehensive comparisons. Future work should focus on developing adaptive mechanisms that can dynamically adjust the evaluation scope based on available computational resources and required accuracy levels.
Dependence on authentic user queries and handling of unanswerable queries: SEARA’s effectiveness is inherently tied to the availability and quality of authentic user query datasets, which may not always be accessible or representative. The method assumes that user queries have retrievable answers within the knowledge base, but real-world scenarios often include unanswerable queries - those for which no relevant information exists. Currently, our framework does not distinguish between answerable and unanswerable queries, which may lead to suboptimal evaluation outcomes and wasted computational resources. This limitation stems from the challenge of automatically detecting query answerability without access to ground truth annotations. Developing robust mechanisms to identify and appropriately handle unanswerable queries represents a promising direction for future work, potentially through query difficulty estimation or retrieval confidence scoring techniques.